OpenStudy (shinebright1):

Which of the following statements is false? The product of two rational numbers is always irrational. The sum of a rational number and an irrational number is always irrational. The product of a nonzero rational number and an irrational number is always irrational. The product of two irrational numbers is either rational or irrational.

OpenStudy (zarkon):

\[\frac{1}{2}\times\frac{1}{3}=\frac{1}{6}\]

OpenStudy (shinebright1):

@Zarkon what does that mean

OpenStudy (zarkon):

i took two rational numbers and multiplied them together to get another rational number

OpenStudy (zarkon):

that should answer your original question